OpenStudy (anonymous):

Simplify by taking roots of the numerator and denominator. sqrt 16 / 81

OpenStudy (anonymous):

\[\sqrt{\frac{16}{81}}=\frac{\sqrt{16}}{\sqrt{81}}=\frac{4}{9}\]

OpenStudy (anonymous):

that matches my answer thanks

OpenStudy (anonymous):

yw